1. find the decimal equivalent of the mixed number for 7 4/5
2. Subtract -5.6 - (-7)
3. Add -21 + (-7)
Please note that all work must be shown
Thank you very much - final exam due tomorrow 8/14

Found 2 solutions by Alan3354, rfer:
Answer by Alan3354(69443)

1. find the decimal equivalent of the mixed number for 7 4/5
= 7.8
-----------
2. Subtract -5.6 - (-7)
= -5.6 + 7
= 7 - 5.6
= 1.4
---------
3. Add -21 + (-7)
= -21 -7
= -28

Answer by rfer(16322)

1)
7.8
2)
-5.6+7=1.4
3)
=-28